OpenStudy (dino):

if x=8 then 2x+3x=.......................

geerky42 (geerky42):

2x + 3x → 5x 5x = ?

OpenStudy (callisto):

Hello! Please post your question in the mathematics group, thanks! Link to mathematics group - http://openstudy.com/study#/groups/Mathematics

OpenStudy (anonymous):

2(8)+3(8)= 16+24=40 A=40